OpenStudy (anonymous):

How do you evaluate the integral 3/2xe^x dx

OpenStudy (psymon):

Do you know how to do integration by parts?

OpenStudy (anonymous):

Well I learned that in Cal 2 but don't remember lol..

OpenStudy (psymon):

Yeah, that would be where ya learn it, haha. Well, there are two formulas that I know of for it, but this is the one I use: \[u \int\limits_{}^{}v - (\int\limits_{}^{}u'\int\limits_{}^{}v)\]

OpenStudy (psymon):

Weird to type that on here, I swear.

OpenStudy (psymon):

But okay, so we need to choose a u and we need to choose a v. We choose our u based on what is convenient to take the derivative of and we choose our v based on what is convenient to take an integral of. Kinda recall some of that?

OpenStudy (anonymous):

Integration using parts

OpenStudy (anonymous):

Yes its coming back sort of lol

OpenStudy (psymon):

Lol, alright, cool. So if we had to pick a u (easy or convenient to differentiate) and a v (easy or convenient to integrate), what would we choose? :P

OpenStudy (anonymous):

I would choose 3/2x to be u and e^x to be v

OpenStudy (psymon):

You don't need to quite do that. Remember, we can factor out that 3/2 and just multiply it in later. Might be easier to do.

OpenStudy (anonymous):

Also how do you know when to use integration by parts?

OpenStudy (anonymous):

so x would be u and e^x would be v

OpenStudy (psymon):

Usually you have something like xe^x, e^x(sinx), sin(x)ln(x), there are usually terms that do not normally match up in terms of normal algebra. Of course you can also know because u-substitution, x-substitution, or any of that won't work. And be careful, you can't just choose x or e^x, those are in the denominator, so they must be (1/x) = u and (1/e^x) = v. Or if you prefer, x^-1 = u, e^(-x) = v

OpenStudy (anonymous):

Sorry I wrote the question wrong -.-. Its actually 3/2 * xe^x dx . The xe^x are not in the deniminator

OpenStudy (psymon):

Oh!. Okay, that makes this integral wayyyy easier. (nightmare if its in denom). Okay, so yes, u = x and v = e^x. So now v itself isn't really needed, we just need integral v.....which is e^x, haha. So if you got the formula, try plugging in the values and see what ya get. Does the formula make sense?

OpenStudy (anonymous):

Ill try it right now

OpenStudy (psymon):

kk, ill do it, too.

OpenStudy (anonymous):

wait wouldn't e^x be the dv?

OpenStudy (anonymous):

and v ends up being e^x still

OpenStudy (psymon):

Yes, it does stay as e^x after integrating it. That's why we picked v to be e^x. it's integral is easy to do. It doesn;t matter that it's the same exact thing, you still plugit into theformula.

OpenStudy (anonymous):

So I got 3/2 * xe^x - 3/2e^x

OpenStudy (psymon):

+ C xDD

OpenStudy (anonymous):

oh yeah that C...Lol thanks again :D

OpenStudy (psymon):

Haha, yeah, I hate the damn thing, too -_- But yep, np ^_^

OpenStudy (anonymous):

I really should know these things but I forgot a lot of stuff from Cal 2 and im paying for it now in Cal 3 -.-

OpenStudy (psymon):

The only reason I know it is because I'm a math tutor at my college. Itforces me to review a lot. I'm taking discrete math and diff eqs here in a couple weeks. I'd be in calc 3, too, but schedule conflict x_x

OpenStudy (anonymous):

ahhh lol. I took discrete math already I believe. I don't have to take diff eqs thank god. I might have to tutor or something so I can remember this stuff..

OpenStudy (psymon):

How was discrete? It looks like a math logic class, haha. But yeah, it helps me remember for sure. I just had to get a recommendation, take a test, then do an interview. I swear, you get enough people coming in for calculus and you never forget it xD

OpenStudy (anonymous):

Haha nice. Discrete wasn't too bad. yes its basically logic and proofs. Are you a CS major? I had to take discrete only because I was a CS major but when I switched to math it wasn't a requirement.

OpenStudy (psymon):

Nah, im actually a pure math major. I still have to take the class, though. I have that class, diff eqs, and calc 3 left before 300-400 level stuff. I do have to do some CSstuff, though, I'm taking a CS class in the fall as well.

OpenStudy (anonymous):

haha have fun with CS. I like computers and all but I couldn't program for pellet...

OpenStudy (anonymous):

wow pellet lol

OpenStudy (psymon):

Yeah, seems like thats exactly what im in for :/ Textbook itself is C++ programming.

OpenStudy (anonymous):

ah I did Java while I was in CS until they changed the entire thing to start in C++

OpenStudy (psymon):

Its really not my thing, but looks like I have to do a few things that could be related to CS. Forgot, I also have to take matrix algebra x_x

OpenStudy (anonymous):

Ah Im guessing Matrix Algebra is basically Linear Algebra? I just took that last month in a summer semester. You could imagine everything was compressed into a month >.< lol it's not too bad. I prefer linear algebra over calculus 3!

OpenStudy (psymon):

Oh wow, lol. Yeah, there were no math classes I could take in the summer or I would have. And yeah, pretty much is. I eventually have to do linear algebra, too. Im going to try and do that in the same semester as calc 3. Is calc 3 really that bad or is just linear algebra more fun? xD

OpenStudy (anonymous):

both..lol I guess you could say Calculus 3 is harder if you don't remember stuff from the previous cal's.. But calculus 2 was definitely worst than calculus 3.

OpenStudy (anonymous):

you'll do fine in cal 3 though because you've been helping me with calculus 3 stuff..lol

OpenStudy (psymon):

I think I just lucked out for calc 2. We were allowed formula sheets, made the class very easy. Now when I took calc 1, guy couldnt teach worth a damn, I was basically self-taught in that class. And yeah, someof the calc 3 stuff isnt bad xD Ive perused through the topics and picked up on some of the ones I thought were simpler. So i've seen the partial derivatives and double/triple integrals. Not much more than that, though.

OpenStudy (anonymous):

Nice. Im basically teaching myself for calculus 3 right now lol. Yeah partial derivatives isn't bad. We just covered triple integrals a couple of days ago. Still doing the hw for double integrals -.- So much hw....

OpenStudy (psymon):

Not good :/ How far into the class are you? Like, how much more do you have to worry about learning?

OpenStudy (anonymous):

I think about 2 weeks left O.O.. thats the scary thing lol. I have my 2nd midterm on Monday and then I think the final is on Friday..

OpenStudy (anonymous):

This is also a 1 month summer course

OpenStudy (psymon):

Ah. Haha, speed calc 3 course, geez O.o Sounds bad. Well, the only reason I ever found this site was because of this paul's online notes thing. Lot of people have seen it, lot of peoplel haven't. But anything I've seen from calc 3 I studied there. You know of the site?

OpenStudy (anonymous):

Nope I never heard of it but I just looked it up. Maybe it'll come in handy though :p.

OpenStudy (psymon):

Yeah, professor put his whole class notes on this open website, so that's pretty awesome. Might help out. And if you come up with another question maybe I can refer to it and help out, too, lol.

OpenStudy (anonymous):

haha sweet. yeah ill check it out

OpenStudy (psymon):

Mhm. I'll look through it some, too, I need to learn it anyway xD I'm looking at something that might be upcoming. Double integrals in polar coordinates.

OpenStudy (anonymous):

yep we did that but I haven't taught myself yet lol

OpenStudy (psymon):

Ah, already did it xD Yeah, im just looking through it now. I was wondering when Id see that again. Hasnt been since trig really that I did much with that.

OpenStudy (anonymous):

lol yeah you will deal with polar coordinates in cal 3 >.<

OpenStudy (psymon):

Yeah, just going through it now. This first example problem doesn't look too bad. Just the whole inequality thing is a new concept.

OpenStudy (anonymous):

for the integral of (x*y*e^xy^(2)) how would I tackle that? Using IBP again?

OpenStudy (anonymous):

with respect to y

OpenStudy (psymon):

\[\int\limits_{}^{}xye ^{xy ^{2}}dy\] Is that correct? You only need it in respect to y?

OpenStudy (anonymous):

correct

OpenStudy (psymon):

Then yeah, we just pretend that x is merely a constant and go from there. In this case it comes off nicely. Just let u = xy^2 and pretty much everything works out.

OpenStudy (psymon):

Come up with something?

OpenStudy (anonymous):

ya and its totally off lol.. I left u = xy^2, v= y^2/2, du = 2xy, dv = y Plugging into IBP formula: uv - integral(vdu) => xy^2 * y^2/2 - integral(y^2/2*y) dy

OpenStudy (anonymous):

* let not left

OpenStudy (psymon):

Lol, we dont need IBP, just treat it like a normal integral o.o

OpenStudy (psymon):

I meant u as in u-substitution o.o

OpenStudy (anonymous):

rofl oops >.< brb

OpenStudy (psymon):

Lol, alright.

OpenStudy (psymon):

I'm gonna have to head out, so I'll just show ya most of it before I go: \[\int\limits_{}^{}xye ^{xy ^{2}}dy\] let u = xy^2, meaning du = 2xydy and dy = (du/2xy) \[\int\limits_{}^{}xye ^{u}\frac{ du }{ 2xy }\] From there you can kinda see how it'll go. I just pretended x was a constant and started integrating in respect to y. Should be able to finish there. Later.

OpenStudy (anonymous):

@psymon I see that part you did above but what now that you have u in the integral?

OpenStudy (psymon):

Well, since I called that portion of the integral u, I keep it as u until after I integrate. Basically, you treat it as if it were a simple x up there. As a formula for the integration of e^x, its actually listed as e^u. So if you got everything up to that point out of the way, you would integrate e^u, which is just e^u, multiply the 1/2 back in and replace u with the original xy^2. If that makes any sense.

OpenStudy (anonymous):

ok got it thx

OpenStudy (psymon):

Okay, cool xD glad it made some sense then.